OpenStudy (anonymous):

perimeter=23 feet x ft (x+3)ft find the value for the triangle or rectangle

OpenStudy (anonymous):



OpenStudy (helder_edwin):

if it is a rectangle then \[ \large 23=2x+2(x+3)=4x+6 \] \[ \large 23-6=4x \] \[ \large 17=4x \] \[ \large x=17/4 \]

OpenStudy (anonymous):

where did u get 2x+2??

OpenStudy (anonymous):

and the 4x+6?

OpenStudy (helder_edwin):

\[ \large 2x+2(x+3)=2x+2x+6=4x+6 \]

OpenStudy (anonymous):

whattt?

OpenStudy (anonymous):

i don't get it? :c

OpenStudy (anonymous):

why is it 2x+2

OpenStudy (helder_edwin):

what do u mean by "2x+2", there is no such thing in my post i wrote \[ \large 2x+2(x+3) \]

OpenStudy (anonymous):

yes that where do u find the formula to solve it? like how?

OpenStudy (helder_edwin):

the perimeter of a rectangle is the addition of the lengths of its four sides, which are equal two-by-two.

OpenStudy (helder_edwin):